# Gromov-Hausdorff convergence of metric spaces of UCP maps

### v1

@@ Abstract.
Acknowledgement: The authors are thankful to Walter van Suijlekom whose India visit under the Jubilee Chair Professorship of the Indian of Academy of Sciences was pivotal for this paper.
@@ 1. Introduction
Given an operator system $E$ and a (complex, separable) Hilbert space $\mathcal{K}$, let $\mathcal{UCP}_{\mathcal{K}}(E)$ denote the collection of all unital completely positive (UCP) maps from $E$ into the algebra of all bounded operators $\mathcal{B}(\mathcal{K})$. We shall show in this note that the general conditions due to van Suijlekom mentioned above actually guarantee much more, viz., the Gromov-Hausdorff convergence of sequences of sets of UCP maps. In particular, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ both converge in the Gromov-Hausdorff metric to $\mathcal{UCP}_{\mathcal{K}}(C(S^{1}))$ as $n$ tends to infinity, implying that for large $n$, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ are close in the Gromov-Hausdorff metric.
The main result is Theorem 1 where we consider a sequence of operator system spectral triples $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ related to an operator system spectral triple $(\mathcal{E},\mathcal{H},D)$ by a $\mathrm{C}^{1}$-approximate complete order isomorphism. Section 4 contains applications of the theorem to examples like the ones mentioned above. The $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ are often called spectral truncations of $(\mathcal{E},\mathcal{H},D)$. Spectral truncations have a widespread appeal - see [4], [8], [13] and [11] for noncommutative geometry, [7] for compact quantum groups, and [5] for machine learning.
@@ 2.1. The BW-topology
For Banach spaces $X$ and $Y$, let $\mathcal{B}(X,Y^{*})$ denote the bounded linear maps of $X$ into $Y^{*}$. For fixed vectors $x$ in $X$ and $y$ in $Y$, define a linear functional $x\otimes y$ on $\mathcal{B}(X,Y^{*})$ by $x\otimes y(L)$ = $L(x)(y)$. Let $Z$ denote the closed linear span in $\mathcal{B}(X,Y^{*})$ of these elementary tensors. Then $\mathcal{B}(X,Y^{*})$ is isometrically isomorphic to $Z^{*}$ with the duality given by
@@
If $Y$ is the algebra $\mathcal{B}(\mathcal{H})$ of bounded operators on a Hilbert space $\mathcal{H}$, then a bounded net $L_{\lambda}$ in $\mathcal{B}(X,\mathcal{B}(H))$ converges in the BW topology to $L$ if and only if $\langle L_{\lambda}(x)h,h^{\prime}\rangle$ converges to $\langle L(x)h,h^{\prime}\rangle$ for all $h$, $h^{\prime}$ in $\mathcal{H}$ and $x$ in $X$. Naturally, $\mathcal{UCP}_{\mathcal{K}}(E)$ inherits the BW-topology from $\mathcal{B}(E,\mathcal{B}(K))$. It is known that $\mathcal{UCP}_{\mathcal{K}}(E)$ is compact in the BW-topology [9, Theorem 7.4].
@@
Let $(\Omega,d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega})$ a compact metric space. Consider the set of all Lipschitz functions
$\text{Lip}(\Omega):=\{f:\Omega\to\mathbb{C}\text{ such that }\|f\|_{1}:=\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{d_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega}(x,y)}<\infty\}.$ (2.1)
@@
Clearly, symmetry and that the triangle inequality hold. Also, the implication $\varphi=\psi\implies d(\varphi,\psi)=0$ is immediate. Suppose now $d(\varphi,\psi)=0$. Then for all $f\in\text{Lip}(\Omega)$ and for all $p$,
$\left\langle(\varphi-\psi)(f)k_{p},k_{p}\right\rangle=0\implies\left\langle(\varphi-\psi)(f)k,k\right\rangle=0$ (2.3)
for all $k\in\mathcal{K}$. For $h,k\in\mathcal{K}$, (2.3) and the polarization identity
@@
together yield $\left\langle(\varphi-\psi)(f)h,k\right\rangle=0$. Hence $(\varphi-\psi)(f)=0$ for all $f\in\text{Lip}(\Omega)$, and consequently $\varphi=\psi$. This completes the proof.
@@
Let $\{\varphi_{\lambda}\}_{\lambda\in\Lambda}$ be a net in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$ such that $\varphi_{\lambda}$ converges to some $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$. Our claim is that $\varphi_{\lambda}$ also converges to $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$. Suppose not. Then there exist $\varepsilon>0$ and an infinite subset $\hat{\Lambda}\subseteq\Lambda$ such that $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for all $\lambda\in\hat{\Lambda}$. Without loss of generality, we may assume that $\hat{\Lambda}=\mathbb{N}$. Thus, we have $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for $\lambda\in\mathbb{N}$. Therefore, there exists a sequence $\{f_{\lambda}\}$ in $\text{Lip}(X)$ with $\|f_{\lambda}\|\leq 1$ and $\|f_{\lambda}\|_{1}\leq 1$ such that
@@
Since $\|f_{\lambda}\|_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}\leq 1$, so $\{f_{\lambda}\}$ is a sequence of equicontinuous functions. Next, we also note that the series in (2.4) is convergent uniformly in $\lambda$ as
$\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|\leq 2.$
Hence, for the given $\varepsilon>0$, there exist a $N_{0}\in\mathbb{N}$ such that
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p-1}}<\varepsilon.$
Consequently, for every $\lambda$
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|<\varepsilon.$ (2.5)
Therefore, (2.4) and (2.5) together give
$\sum_{p=1}^{N_{0}}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|>\varepsilon.$ (2.6)
Since there are only finitely many $p$ in (2.6), we can choose a $p_{0}\in\{1,\ldots,N_{0}\}$ such that
@@ 2.3. A metric for the BW-topology when $E$ is a finite-dimensional operator system
$d_{E}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$ (2.9)
@@
Obviously, in the finite dimensional case, one can omit the condition $\|f\|_{1}\leq 1$ in the definition of the metric. The metric could be defined as $d_{E}^{\prime}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1\right\}$. This too would metrize the BW-topology. So, the two metrics $d_{E}$ and $d_{E}^{\prime}$ are equivalent. In other words, the Dirac operator is not necessary to define the metric. We introduce the Dirac operator at this stage because it is essential in formulating the sufficient conditions which will guarantee convergence in the next section.
@@ 3. The Gromov–Hausdorff convergence
Now that we know about the metric space structure of $\mathcal{UCP}_{\mathcal{K}}(E)$, we turn to the Gromov-Hausdorff metric $d_{GH}(M_{1},M_{2})$ between two compact metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$.
@@
    for all $f\in\mathcal{E}$ and $h\in\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{E}_{n}$.
@@
The following results transfer [13, Proposition 4] to the setting of $\mathrm{ucp}$ maps.
@@
$d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))\leq\frac{1}{2}\operatorname{dis}\mathfrak{R}_{n}.$ (3.4)
@@ 4.1. The Fejér–Riesz operator systems
$\left(\mathcal{E}:=C^{\infty}(S^{1}),\mathcal{H}:=L^{2}(S^{1}),D:=-i\frac{d}{d\theta}\right).$ (4.1)
@@
$\left(\mathcal{E}_{n}:=C(S^{1})_{(n)},\mathcal{H}_{n}:=L^{2}(S^{1}),D_{n}:=-i\frac{d}{d\theta}\right).$ (4.2)
@@
where we recall that $F_{n}=\sum_{|k|\leq n-1}(1-|k|/n)e^{ik\theta}$ is the Fejér kernel (cf. [12, Chapter 2]). Also, define
@@
From the definition and the properties of the kernel $F_{n}$, it follows that $R_{n}$ and $S_{n}$ are both unital, positive, contractive, and Lipschitz contractive. Since $C(S^{1})$ is a commutative $C^{*}$-algebra, by [9, Theorems 3.9 and 3.11], $R_{n}$ and $S_{n}$ are ucp maps. Therefore, $(R_{n},S_{n})$ form a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$. Hence, we can apply Theorem 1 to these operator systems.
@@ 4.3. The $d$-dimensional torus
Here, $\mathcal{E}$ represents the *-algebra of smooth functions on the torus $\Omega$. These functions act by multiplication on a Hilbert space, denoted by $\mathcal{H}$, which consists of $\mathrm{L}^{2}$-sections of the spinor bundle $\textit{spb}(\Omega)$. We identify $\textit{spb}(\Omega)$ with the trivial bundle $\Omega\otimes V$, where $V:=\mathbb{C}^{\lfloor\frac{d}{2}\rfloor}$, and we write $\mathcal{H}:=\mathrm{L}^{2}(\Omega)\otimes V\cong\mathrm{L}^{2}(\textit{spb}(\Omega))$ for the Hilbert space. The Dirac operator $D$ given by
$D=-i\sum_{\mu=1}^{\lfloor\frac{d}{2}\rfloor}\partial_{\mu}\otimes\gamma^{\mu}$
@@
For an integer $N\geq 0$, let $P_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ of the eigenvalues $\lambda$ with $|\lambda|\leq N$. More concretely, we have $P_{N}\mathcal{H}=\mathrm{span}\{e_{n}\,:\,n\in\mathbb{Z}^{d},\|n\|\leq N\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.

### v2

@@ 1. Introduction
Given a closed operator system $E$ and a (complex, separable) Hilbert space $\mathcal{K}$, let $\mathcal{UCP}_{\mathcal{K}}(E)$ denote the collection of all unital completely positive (UCP) maps from $E$ into the algebra of all bounded operators $\mathcal{B}(\mathcal{K})$. We shall show in this note that the general conditions due to van Suijlekom mentioned above actually guarantee much more, viz., the Gromov-Hausdorff convergence of sequences of sets of UCP maps. In particular, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ both converge in the Gromov-Hausdorff metric to $\mathcal{UCP}_{\mathcal{K}}(C(S^{1}))$ as $n$ tends to infinity, implying that for large $n$, $\mathcal{UCP}_{\mathcal{K}}(C(S^{1})_{(n)})$ and $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{n})$ are close in the Gromov-Hausdorff metric.
The main result is Theorem 1 where we consider a sequence of operator system spectral triples $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ related to an operator system spectral triple $(\mathcal{E},\mathcal{H},D)$ by a $\mathrm{C}^{1}$-approximate complete order isomorphism. Section 4 contains applications of the theorem to examples like the ones mentioned above as well as for polyhedral truncations motivated by works of Travaglini [17]. The $\{(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})\}_{n}$ are often called spectral truncations of $(\mathcal{E},\mathcal{H},D)$. Spectral truncations have a widespread appeal - see [4], [11], [16] and [14] for noncommutative geometry, [10] for compact quantum groups, and [5] for machine learning.
@@ 2.1. The BW-topology
For Banach spaces $X$ and $Y$, let $\mathcal{B}(X,Y^{*})$ denote the bounded linear maps of $X$ into $Y^{*}$. For fixed vectors $x$ in $X$ and $y$ in $Y$, define a linear functional $x\otimes y$ on $\mathcal{B}(X,Y^{*})$ by $x\otimes y(L)$ = $L(x)(y)$. Let $Z$ denote the closed linear span in $(\mathcal{B}(X,Y^{*}))^{*}$ of these elementary tensors. Then $\mathcal{B}(X,Y^{*})$ is isometrically isomorphic to $Z^{*}$ with the duality given by
@@
If $Y^{*}$ is the algebra $\mathcal{B}(\mathcal{H})$ of bounded operators on a Hilbert space $\mathcal{H}$, then a bounded net $L_{\lambda}$ in $\mathcal{B}(X,\mathcal{B}(H))$ converges in the BW topology to $L$ if and only if $\langle L_{\lambda}(x)h,h^{\prime}\rangle$ converges to $\langle L(x)h,h^{\prime}\rangle$ for all $h$, $h^{\prime}$ in $\mathcal{H}$ and $x$ in $X$. Naturally, $\mathcal{UCP}_{\mathcal{K}}(E)$ inherits the BW-topology from $\mathcal{B}(E,\mathcal{B}(K))$. It is known that $\mathcal{UCP}_{\mathcal{K}}(E)$ is compact in the BW-topology [12, Theorem 7.4].
@@
Let $(\Omega,d_{\Omega})$ be a compact metric space. Consider the set of all Lipschitz functions
$\text{Lip}(\Omega):=\{f:\Omega\to\mathbb{C}\text{ such that }\|f\|_{1}:=\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{d_{\Omega}(x,y)}<\infty\}.$ (2.1)
@@
Clearly, symmetry and the triangle inequality hold. Also, the implication $\varphi=\psi\implies d(\varphi,\psi)=0$ is immediate. Suppose now that $d(\varphi,\psi)=0$. Then for all $f\in\text{Lip}(\Omega)$ and for all $p$, we have $\left\langle(\varphi-\psi)(f)k_{p},k_{p}\right\rangle=0$. By denseness, this implies that
$\left\langle(\varphi-\psi)(f)k,k\right\rangle=0\text{ for all }k\in\mathcal{K}.$ (2.3)
For $h,k\in\mathcal{K}$, (2.3) and the polarization identity
@@
together yield $\left\langle(\varphi-\psi)(f)h,k\right\rangle=0$. Hence $(\varphi-\psi)(f)=0$ for all $f\in\text{Lip}(\Omega)$, and consequently $\varphi=\psi$. This completes the proof. ∎
The following is probably known. We give a proof because we could not find a suitable reference in the literature.
@@
Let $\{\varphi_{\lambda}\}_{\lambda\in\Lambda}$ be a net in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$ such that $\varphi_{\lambda}$ converges to some $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),BW)$. Our claim is that $\varphi_{\lambda}$ also converges to $\varphi$ in $(\mathcal{UCP}_{\mathcal{K}}(C(\Omega)),d)$. Suppose not. Then there exist $\varepsilon>0$ and an infinite subset $\hat{\Lambda}\subseteq\Lambda$ such that $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for all $\lambda\in\hat{\Lambda}$. Invoking the Axiom of Choice, we may assume that $\hat{\Lambda}=\mathbb{N}$. Thus, we have $d(\varphi_{\lambda},\varphi)>2\varepsilon$ for $\lambda\in\mathbb{N}$. Therefore, there exists a sequence $\{f_{\lambda}\}$ in $\text{Lip}(X)$ with $\|f_{\lambda}\|\leq 1$ and $\|f_{\lambda}\|_{1}\leq 1$ such that
@@
Since $\|f_{\lambda}\|_{1}\leq 1$, so $\{f_{\lambda}\}$ is a sequence of equicontinuous functions. Next, in view of
$\left|\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\rangle\right|\leq 2$
and the fact that for a given $\varepsilon>0$, there exist an $N_{0}\in\mathbb{N}$ such that
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p-1}}<\varepsilon,$
we have
$\sum_{p=N_{0}+1}^{\infty}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|<\varepsilon$ (2.5)
for every $\lambda$. Therefore, (2.4) and (2.5) together give
$\sum_{p=1}^{N_{0}}\frac{1}{2^{p}}\left|\left\langle(\varphi_{\lambda}-\varphi)(f_{\lambda})k_{p},k_{p}\right\rangle\right|>\varepsilon\;\;\text{ for every $\lambda$}.$ (2.6)
Since there are only finitely many $p$ in (2.6), we can choose a $p_{0}\in\{1,\ldots,N_{0}\}$ such that after passing to a subsequence and continuing to denote the subsequence by $\lambda$, we have
@@ 2.3. A metric for the BW-topology when $E$ is a finite-dimensional operator system
$d_{E}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1,\|f\|_{1}\leq 1\right\}$
@@
Obviously, in the finite dimensional case, one can omit the condition $\|f\|_{1}\leq 1$ in the definition of the metric. The metric could be defined as
$d_{E}^{\prime}(\varphi,\psi)=\sup\left\{\sum_{p=1}^{\infty}\frac{1}{2^{p}}\left|\langle(\varphi-\psi)(f)k_{p},k_{p}\rangle\right|:f\in\mathcal{E},\|f\|\leq 1\right\}.$
This too would metrize the BW-topology. So, the two metrics $d_{E}$ and $d_{E}^{\prime}$ are equivalent. In other words, the Dirac operator is not necessary to define the metric. We introduce the Dirac operator at this stage because it is essential in formulating the sufficient conditions which will guarantee convergence in the next section.
@@ 3. The Gromov–Hausdorff convergence
In this section, we shall consider the usual Gromov-Hausdorff metric $d_{GH}(M_{1},M_{2})$ between two compact metric spaces $(M_{1},d_{1})$ and $(M_{2},d_{2})$ and apply it to measure distances between operator system spectral triples. For various notions of distances between Lip-normed operator systems and inter-relations between them, we refer the reader to section 3 of [9].
@@
    for all $f\in\mathcal{E}$ and $h\in\mathcal{E}_{n}$.
@@
The following results transfer [16, Proposition 4] to the setting of $\mathrm{ucp}$ maps.
@@
$d_{GH}(\mathcal{UCP}_{\mathcal{K}}(E_{n}),\mathcal{UCP}_{\mathcal{K}}(E))\leq\frac{1}{2}\operatorname{dis}\mathfrak{R}_{n}.$
@@
A different proof for finite-dimensional $\mathcal{K}$ can be given using [10, Proposition 5.19] of Leimbach. This was communicated to us by him. He showed a stronger result, viz., an estimate on the complete Gromov-Hausdorff distance ([8, Definition 3.2]). The idea of his proof stems from Proposition 2.14 of [7] which uses the notion of bridge introduced by Rieffel [13].
@@ 4.1. The Fejér–Riesz operator systems
$\left(\mathcal{E}:=C^{\infty}(S^{1}),\mathcal{H}:=L^{2}(S^{1}),D:=-i\frac{d}{d\theta}\right).$
@@
$\left(\mathcal{E}_{n}:=C(S^{1})_{(n)},\mathcal{H}_{n}:=L^{2}(S^{1}),D_{n}:=-i\frac{d}{d\theta}\right).$
That these two spectral triples have the BW property is a consequence of results of section 2.
@@
where we recall that $F_{n}=\sum_{|k|\leq n-1}(1-|k|/n)e^{ik\theta}$ is the Fejér kernel (cf. [15, Chapter 2]). Also, define
@@
From the definition and the properties of the kernel $F_{n}$, it follows that $R_{n}$ and $S_{n}$ are both unital, positive, contractive, and Lipschitz contractive. Since $C(S^{1})$ is a commutative $C^{*}$-algebra, by [12, Theorems 3.9 and 3.11], $R_{n}$ and $S_{n}$ are ucp maps. Therefore, $(R_{n},S_{n})$ form a $C^{1}$-approximate complete order isomorphism for $(\mathcal{E}_{n},\mathcal{H}_{n},D_{n})$ and $(\mathcal{E},\mathcal{H},D)$. Hence, we can apply Theorem 1 to these operator systems.
@@ 4.3. The $d$-dimensional torus
Here, $\mathcal{E}$ represents the *-algebra of smooth functions on the torus $\Omega$. These functions act by multiplication on a Hilbert space, denoted by $\mathcal{H}$, which consists of $\mathrm{L}^{2}$-sections of the spinor bundle $\textit{spb}(\Omega)$. We identify $\textit{spb}(\Omega)$ with the trivial bundle $\Omega\otimes V$, where $V:=\mathbb{C}^{d}$, and we write $\mathcal{H}:=\mathrm{L}^{2}(\Omega)\otimes V\cong\mathrm{L}^{2}(\textit{spb}(\Omega))$ for the Hilbert space. The Dirac operator $D$ given by
$D=-i\sum_{\mu=1}^{d}\partial_{\mu}\otimes\gamma^{\mu}$
@@
4.3.1. Spherical truncation
For an integer $N\geq 0$, let $P_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ with $|\lambda|\leq N$. More concretely, we have $P_{N}\mathcal{H}=\mathrm{span}\{e_{n}\,:\,n\in\mathbb{Z}^{d},\|n\|\leq N\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.